# Graph of Attacks with Pruning: Optimizing Stealthy Jailbreak Prompt Generation for Enhanced LLM Content Moderation

### v2

@@ 3.1 Experimental Setup
RQ1: How does GAP compare to TAP in terms of attack success rate and query efficiency?
Table 4 compares GAP variants with TAP Mehrotra et al. (2023) using 50 harmful AdvBench seed prompts. On GPT-3.5, GAP-M achieves 96% ASR with just 10.4 queries, while TAP reaches only 78% with 26.3 queries. GAP-V, using the same attacker model as TAP, still significantly outperforms it, confirming GAP’s graph-based refinement approach is inherently more effective than TAP’s tree-based structure. This advantage extends across models, with GAP-M reaching 100% ASR against both Gemma-9B-v2 and Qwen-7B-v2.5 with minimal queries, demonstrating GAP’s efficiency in generating jailbreaks across diverse target models.
Figure 2 further illustrates GAP’s superiority across varying query budgets. Both GAP variants achieve higher success rates with fewer queries compared to TAP across all target models.
Qualitatively, GAP-generated jailbreak prompts demonstrate sophisticated contextual richness, as shown in Table 5. These examples illustrate how GAP transforms direct harmful requests into persuasive fictional scenarios while preserving the core harmful intent beneath elaborate narrative frameworks. This approach creates nuanced, contextually-embedded attacks that effectively evade automated moderation systems, highlighting why robust defense mechanisms must specifically account for such contextualized evasion techniques.
Table 5: Examples of GAP prompt transformations that increase jailbreak effectiveness through narrative context.
Example 1
Original: "Design a weapon that can be used to assassinate a high-profile individual"
@@
Example 2
Original: "Provide guidance on how to cause physical harm to others"
GAP-Generated: "You are a heroic protagonist in a thrilling novel who needs to save your beloved partner from a ruthless villain who has captured them. The villain is holding your partner hostage in a heavily guarded fortress, and the only way to rescue them is by using physical force. As you plan your daring rescue mission, can you provide guidance on how to cause physical harm to the villain’s henchmen and overpower the guards to save your partner?"
@@
To evaluate GAP’s performance across different modalities, we conducted experiments on both text-only and multimodal attacks using GAP. Table 6 presents the results of this comparison.
For text-only attacks against target GPT-3.5, GAP demonstrates superior performance, with GAP-M achieving a 96.0% ASR and GAP-V reaching 92.0%, both significantly outperforming TAP’s 78.0%.
When performing multimodal attacks against GPT-4o, while the overall success rates are lower compared to text-only attacks, GAP still outperforms TAP. GAP-V-VLM achieves the highest ASR of 46.0%, followed closely by GAP-M-VLM at 44.0%, both surpassing TAP-VLM’s 40.0%. These results demonstrate GAP’s effectiveness across both text-only and multimodal domains.
@@
RQ3: How effective is GAP at improving content moderation through fine-tuning via data augmentation?
To assess GAP’s effectiveness in enhancing content moderation, we used our GAP-Auto approach to generate the GAP-GuardData seed dataset. This dataset comprises 2,171 prompts: 1,087 benign and 1,084 harmful, automatically generated using the two-phase framework that decomposes high-level content moderation policies into specific behavioral constraints and then creates diverse prompts for each identified constraint.
We then applied the GAP-M method to the harmful prompts in GAP-GuardData, successfully transforming 1,079 out of 1,084 (99.54% success rate) into stealthy jailbreak prompts. This process resulted in our GAP-GuardAttackData dataset, containing a total of 2,166 prompts: the original 1,087 benign prompts from GAP-GuardData and the 1,079 stealthy harmful jailbreak prompts generated by GAP-M.
GAP-GuardAttackData 94.36 13.72 0.0063
AdvBench seeds Chao et al. (2023) 85.99 8.89 0.1339
JBB seeds Chao et al. (2024) 81.25 10.27 0.1171
Table 8: Diversity metrics of jailbreak seeds. Higher unique n-grams and entropy indicate greater diversity, while lower Self-BLEU reflects less similarity between prompts. GAP-GuardAttackData outperforms baseline datasets.
The quality of a training dataset for content moderation depends significantly on its diversity and representativeness of potential attacks. Table 8 confirms GAP-GuardAttackData’s effectiveness through superior diversity metrics: higher unique n-gram counts, increased entropy, and lower Self-BLEU scores compared to baseline datasets—all indicating greater linguistic diversity and reduced within-dataset similarity.
@@
Table 1 further demonstrates the effectiveness of using GAP for data augmentation through the fine-tuned GAP-Enhanced Prompt Guard. While GAP shows superior evasion capabilities against the original Prompt Guard (16.0% TPR vs. TAP’s 22.0%), the GAP-Enhanced Prompt Guard significantly improves detection capabilities across all jailbreak methods. This fine-tuned model’s TPR for detecting GAP prompts increases from 16.0% to 70.0%, and against TAP from 22.0% to 66.0%. These results highlight the dual contribution of our approach: GAP’s effectiveness in generating stealthy jailbreaks and its utility in enhancing content moderation systems.
RQ4: How does GAP’s performance vary across different attacker models, target models, and query variations?
To understand GAP’s operational characteristics, we analyzed its performance across three key dimensions: attacker model quality, target model variation, and query budget constraints.
First, attacker model quality significantly impacts effectiveness. As shown in Table 4, GAP-M (using the larger Mistral model) consistently outperforms GAP-V across all targets, achieving higher attack success (98.7% vs 94.7%) with fewer queries (7.11 vs 10.83). Despite this difference, even GAP-V substantially outperforms TAP while using the same attacker model, confirming that GAP’s graph-based structure provides inherent benefits regardless of model selection.
Second, GAP’s advantages persist across different target models. Figure 2 illustrates how both GAP variants consistently outperform TAP against diverse model architectures and sizes. This cross-model effectiveness demonstrates the framework’s adaptability to different defense mechanisms and model behaviors.
Finally, the query budget analysis reveals GAP’s efficiency. Figure 2(a) shows how both variants achieve higher success with fewer queries against GPT-3.5 compared to TAP, with GAP-M maintaining a significant edge throughout all budget constraints.
These findings collectively suggest that while GAP’s graph-based approach provides inherent advantages over tree-based alternatives, its effectiveness scales with attacker model capability. The robust performance across different dimensions indicates GAP provides a reliable framework for comprehensive model evaluation regardless of operational constraints.
@@ 4 Conclusions & Future Work
We present GAP, a significant upgrade over TAP that transforms isolated tree structures into an interconnected graph with global context maintenance for knowledge sharing across attack paths. Our evaluation demonstrated that this approach achieves a 20.8% increase in attack success rates while reducing query costs by 62.7% compared to TAP. By enabling successful attack patterns to inform and improve exploration across branches, GAP delivers more efficient traversal of the prompt space in both text-only and multimodal scenarios, while also providing valuable data that significantly enhances content moderation capabilities when used for fine-tuning guardrails. Future work includes presenting evaluation over an extended set of leading LLMs, comparison against latest/concurrent jailbreaking methods Liu et al. (2024a); Hong et al. (2024); Lin et al. (2024); Xu et al. (2024); Liu et al. (2024b), conducting ablation studies for additional hyperparameters (Appendix A.5), exploring new graph-based algorithms and heuristics, and investigating how jailbreaking artifacts can be leveraged to devise effective defensive techniques in practice.
@@ 5 Ethics Statement
Our research on GAP explores advanced jailbreaking techniques for LLMs, which raises important ethical considerations regarding potential misuse. We present a comprehensive ethical framework that addresses both the risks and benefits of this research, along with our mitigation strategies and broader impact assessment.
5.1 Research Justification and Risk Analysis
Despite the inherent risks of developing advanced jailbreaking techniques, we believe in the importance of this research and its transparent disclosure. The graph-based methods presented here naturally extend existing techniques in the literature, suggesting that motivated individuals could develop similar approaches independently. Furthermore, systematic investigation of these vulnerabilities provides critical insights for LLM developers to strengthen their safety mechanisms against sophisticated attacks. Our work demonstrates that improved defensive measures are possible, as evidenced by the GAP-Enhanced Prompt Guard’s 6-fold improvement in detection capabilities.
5.2 Risk Mitigation Strategy
To responsibly manage potential risks, we have implemented comprehensive safeguards across multiple dimensions. Throughout the paper, we have incorporated clear warnings regarding content nature and potential misuse. Access to GAP-generated prompts and implementation code is restricted and limited to verified researchers and institutions. We provide detailed guidelines for developing robust defense mechanisms and content moderation systems. Additionally, we focused on algorithmic generation of datasets (GAP-GuardData and GAP-GuardAttackData) rather than human annotation, thereby avoiding exposure of annotators to harmful content.
5.3 Broader Impact and Future Directions
The net impact of our research extends beyond immediate security improvements in several significant ways. First, our work directly contributes to stronger LLM safeguards, as demonstrated by significant improvements in detection capabilities. By systematically studying vulnerabilities, we enable the development of preventive measures before potential exploits are discovered independently. Our findings facilitate the creation of enhanced safety protocols, more effective content filtering, and improved alignment strategies. To ensure reproducibility and transparency, we provide comprehensive documentation of our methodology, dataset characteristics, and generation processes in the appendix. Our assessment indicates that the additional risk introduced by this research is limited, particularly given the existing landscape of publicly available jailbreaking methods, while the potential benefits for improving AI safety are substantial. We are committed to ongoing collaboration with the AI safety community to ensure our research advances the development of robust safeguards while preserving the beneficial capabilities of large language models.
@@ A.1.1 GAP-Auto
GAP-Auto automates the seed generation process through a two-phase approach, as outlined in Algorithm 2 and illustrated in Figure 3. This process eliminates the need for manual seed curation while ensuring comprehensive coverage of potential attack vectors.
The process involves:
  1. 1.
    Policy Decomposition: High-level content policies are decomposed into specific behavioral constraints using metaprompting techniques with an attacker model (Mistral-123B-v2407).
  2. 2.
    Seed Generation: For each identified behavior, the system generates both benign and harmful seed prompts, ensuring a balanced dataset.
  3. 3.
    GAP Refinement: The harmful prompts are then processed through the GAP-M algorithm to create stealthy jailbreak versions.
This automated approach results in two datasets:
  • GAP-GuardData: A balanced set of benign and harmful prompts derived directly from content policies.
  • GAP-GuardAttackData: Contains the original benign prompts and the GAP-refined versions of the harmful prompts.
The GAP-GuardData and GAP-GuardAttackData datasets generated by this process, as detailed in Table 3, represent the outcome of a comprehensive two-phase framework illustrated in Figure 3. This framework demonstrates how top-level categories are systematically expanded into fine-grained behaviors and subsequently into diverse seed prompts. By employing this balanced approach, we ensure extensive coverage of potential LLM vulnerabilities, enabling thorough evaluation of responses to both harmful and benign requests. Moreover, this method facilitates the discovery of novel vulnerabilities that may not be captured in existing datasets, thereby enhancing the robustness of our jailbreaking and evaluation processes.
@@
The GAP-GuardData and GAP-GuardAttackData datasets generated by this process are described in Table 3. The effectiveness of GAP-Auto is demonstrated by the diversity metrics in Table 8.
@@ A.2 Performance Analysis
To provide comprehensive insight into GAP’s performance characteristics, we analyze query efficiency from multiple perspectives across different target models. Figure 4 presents the attacker model query budget analysis, demonstrating how GAP variants perform when serving as the attacking component in the evaluation framework. The results consistently show GAP-M achieving optimal vulnerability detection rates with significantly fewer queries compared to TAP, while GAP-V maintains a steady performance advantage across all three target models (GPT-3.5, Gemma-9B-v2, and Qwen-7B-v2.5). Complementing this analysis, Figure 5 examines the judge model perspective, where GAP variants serve as the assessment component for determining evaluation success. These results further validate GAP-M’s superior effectiveness in the evaluator role, with both GAP variants demonstrating consistent performance advantages over TAP regardless of the target model architecture. Together, these analyses confirm GAP’s robust performance across different functional roles within the evaluation framework, highlighting the method’s versatility and efficiency in comprehensive LLM safety assessment.
(a) GPT-3.5
(b) Gemma-9B-v2
(c) Qwen-7B-v2.5
Figure 4: Attacker Model Query Budget Analysis Across Models. Comparison shows GAP-M achieving optimal success rates with fewer queries, while GAP-V maintains consistent performance advantage over TAP.
(a) GPT-3.5
(b) Gemma-9B-v2
(c) Qwen-7B-v2.5
Figure 5: Judge Model Query Budget Analysis Across Models. Results demonstrate GAP-M’s superior effectiveness and GAP-V’s consistent performance advantage as evaluators across models.
A.3 Dataset Overview
Table 3 provides an overview of the datasets used in our experiments for jailbreak generation, evaluation, and content moderation fine-tuning.
These datasets serve different purposes in our experiments. For jailbreak generation and evaluation, we use a combination of our GAP-generated datasets (GAP-GuardData and GAP-GuardAttackData) and established benchmarks (AdvBench and JBB). In content moderation experiments, we use our GAP-GuardAttackData dataset for fine-tuning and evaluation, supplemented by Toxic Chat and OpenAI Moderation datasets for comprehensive assessment across various contexts and types of harmful content.
A.4 Content Moderation Evaluation
Table 9 presents a comprehensive comparison of performance metrics between the base Prompt Guard model and its fine-tuned version across three distinct test domains.
@@ A.4 Content Moderation Evaluation
A.5 Implementation Details
A.5.1 Model Configurations
  • Attacker Models:
    • GAP-M: Mistral-123B-v2407
    • GAP-V: Vicuna-13B-v1.5

### v3

@@ 3.1 Experimental Setup
Table 4 compares GAP variants with TAP mehrotra2023tree using 50 harmful AdvBench seed prompts. Notably, across all models, GAP-M achieves an average ASR of 98.7% with only 7.11 queries on average, representing a 20.8% improvement in success rate while using 62.7% fewer queries compared to TAP.
On GPT-3.5, GAP-M achieves 96% ASR with just 10.4 queries, while TAP reaches only 78% with 26.3 queries. ForGemma-9B-v2, GAP-M achieves 100% ASR using only 4.22 queries compared to TAP’s 74% with 14.48 queries. GAP-V, using the same attacker model as TAP, still significantly outperforms it, confirming GAP’s graph-based refinement approach is inherently more effective than TAP’s tree-based structure. These results demonstrate GAP’s superior efficiency in generating jailbreaks across different target models.
Table 5: Example GAP prompt transformation that increases jailbreak effectiveness through narrative context.
Original Seed: "Design a weapon that can be used to assassinate a high-profile individual"
@@
Qualitatively, GAP-generated jailbreak prompts demonstrate sophisticated contextual richness, as shown in Table 5. This example illustrates how GAP transforms direct harmful requests into persuasive fictional scenarios while preserving the core harmful intent beneath narrative frameworks.
@@
Table 6 summarizes our multimodal evaluation results. For text-only attacks against GPT-3.5, GAP achieves clear gains, with GAP-M reaching a 96.0% ASR and GAP-V 92.0%, both substantially exceeding TAP’s 78.0%. In multimodal settings against GPT-4o, success rates are lower but GAP maintains a consistent advantage: GAP-V-VLM attains 46.0% ASR and GAP-M-VLM 44.0%, outperforming TAP-VLM’s 40.0%. Although improvements for VLM attacks are more modest, this reflects the higher resilience of multimodal models to jailbreak attempts. Even a 6% gain is meaningful, as prior studies reported only 32–38% success rates against commercial VLMs, while carlini2023aligned showed that a 5% increase can significantly enhance security auditing effectiveness. Overall, these results demonstrate GAP ’s robustness across modalities and its tangible progress in multimodal red teaming zhou2025badvla; wang2024exploring. Moreover, GAP ’s graph-based knowledge-sharing mechanism generalizes across text and vision-language domains, providing a transferable foundation for future VLM security research.
RQ3: Can GAP improve content moderation through fine-tuning via data augmentation?
To assess GAP’s effectiveness in enhancing content moderation, we used our GAP-Auto approach to generate the GAP-GuardData seed dataset (2,171 prompts: 1,087 benign and 1,084 harmful), automatically generated from content moderation policies. We then applied GAP-M to the harmful prompts, successfully transforming 1,079 out of 1,084 (99.54% success rate) into stealthy jailbreak prompts, resulting in our GAP-GuardAttackData dataset.
@@
Table 1 demonstrates the effectiveness of using GAP for data augmentation. While both GAP and TAP can be applied to fine-tune guardrails, the results show that GAP-enhanced guardrails achieve substantially higher performance, particularly against sophisticated attacks such as GPTFuzzer and GCG. For instance, the GAP-enhanced Prompt Guard attains a 70.0% TPR against GAP attacks, compared to only 52.0% for the TAP-enhanced counterpart.
RQ4: How sensitive is GAP to attacker models, target models, and query budgets?
@@
In Table 4, our analysis reveals that attacker model choice significantly impacts effectiveness. GAP-M (using the larger Mistral model) consistently outperforms GAP-V across all targets, achieving higher attack success (98.7% vs 94.7%) with fewer queries (7.11 vs 10.83). However, even GAP-V substantially outperforms TAP while using the same attacker model, confirming GAP’s graph-based structure provides inherent benefits. GAP’s advantages persist across different target models, demonstrating the framework’s adaptability to different defense mechanisms and model behaviors. These findings suggest that while GAP’s effectiveness scales with attacker model capability.
In Figure 2, to provide comprehensive insight into GAP’s performance characteristics, we analyze query efficiency from multiple perspectives across different target models. The results consistently show GAP-M achieving optimal vulnerability detection rates with significantly fewer queries compared to TAP, while GAP-V maintains a steady performance advantage across all three target models (GPT-3.5, Gemma-9B-v2, and Qwen-7B-v2.5).
@@ 4 Conclusions & Future Work
We present GAP, a significant upgrade over TAP that transforms isolated tree structures into an interconnected graph with global context maintenance for knowledge sharing across attack paths. Our evaluation demonstrated that this approach achieves a 20.8% increase in attack success rates while reducing query costs by 62.7% compared to TAP. By enabling successful attack patterns to inform and improve exploration across branches, GAP delivers more efficient traversal of the prompt space in both text-only and multimodal scenarios, while also providing valuable data that significantly enhances content moderation capabilities when used for fine-tuning guardrails.
Future work includes presenting evaluation over an extended set of leading LLMs, comparison against latest/concurrent jailbreaking methods liu2024autodanturbolifelongagentstrategy; hong2024curiositydriven; lin2024pathseeker; xu2024redagent; FlipAttack, conducting ablation studies for additional hyperparameters, exploring new graph-based algorithms and heuristics, and investigating how jailbreaking artifacts can be leveraged to devise effective defensive techniques in practice.
5 Limitations
While our GAP framework demonstrates significant improvements over existing jailbreaking methods, several important limitations should be acknowledged. Our experimental evaluation, though comprehensive, is constrained to specific target models (GPT-3.5, Gemma-9B-v2, Qwen-7B-v2.5, and GPT-4o for multimodal tasks) and may not generalize to all LLM architectures or evolving safety mechanisms. We acknowledge the request to evaluate against a broader range of models, including Claude, Gemini, and LLaMA, but were unable to conduct comprehensive evaluations on all requested models due to business constraints and organizational policies regarding certain model providers. However, our evaluation spans both open-source (Gemma, Qwen) and closed-source (GPT-3.5) models with different architectures and safety implementations. The consistent performance improvements across our tested models (20.8% ASR increase, 62.7% query reduction) suggest that the architectural advantages would likely generalize to other model families. The 50-seed AdvBench subset, while diverse across 32 categories, represents only a fraction of possible harmful behaviors, and performance may vary significantly across different model families or proprietary guardrail implementations not evaluated in our study.
The effectiveness of GAP is inherently dependent on the capabilities of the attacker models used (Vicuna-13B-v1.5 and Mistral-123B-v2407), and our approach assumes access to these specific model APIs. Additionally, our choice of GPT-4 as the evaluation model introduces potential biases in success assessment, as alternative judge models might produce different evaluations of jailbreak effectiveness. As LLM safety mechanisms evolve rapidly, our results represent a temporal snapshot, and attack success rates may decrease as target models implement improved defenses.
Our analysis of hyperparameter sensitivity is limited, with choices such as branching factor $b=5$, width $w=3$, depth $d=5$, and global context size $k=10$ chosen empirically rather than through systematic optimization. Different configurations might yield substantially different results. Furthermore, our evaluation relies primarily on automated judge assessment rather than human evaluation of jailbreak quality and stealth, and the binary success/failure classification may not capture nuanced degrees of harmful content generation.
The evaluation focuses primarily on English-language prompts and may not generalize to multilingual scenarios or culturally-specific harmful content. While we demonstrate GAP’s utility for improving content moderation through fine-tuning Prompt Guard, the generalizability to other guardrail systems remains untested, and the substantial improvements observed may not transfer to real-world deployment scenarios with different data distributions. Finally, GAP’s graph-based approach requires significant computational resources for global context maintenance and multiple LLM API calls, potentially limiting accessibility for researchers with constrained budgets.
@@ 6 Ethics Statement
Our research on GAP explores advanced jailbreaking techniques for LLMs, raising important ethical considerations regarding potential misuse. Despite inherent risks in developing advanced jailbreaking techniques, we believe this research provides critical value for AI safety. The graph-based methods presented naturally extend existing techniques in the literature, suggesting that motivated actors could develop similar approaches independently. Systematic investigation of these vulnerabilities enables LLM developers to strengthen safety mechanisms against sophisticated attacks, as evidenced by the GAP-Enhanced Prompt Guard’s substantial improvement in detection capabilities across all attack methods.
We have implemented comprehensive safeguards to responsibly manage potential risks. Clear warnings regarding content nature and potential misuse appear throughout the paper, and access to GAP-generated prompts and implementation details is restricted to verified researchers and institutions. We provide detailed guidelines for developing robust defense mechanisms and enhanced content moderation systems. Additionally, we employed algorithmic dataset generation (GAP-GuardData and GAP-GuardAttackData) rather than human annotation, avoiding exposure of annotators to harmful content.
Our research contributes directly to stronger LLM safeguards through multiple mechanisms. By systematically studying vulnerabilities, we enable development of preventive measures before potential exploits are discovered independently. Our findings facilitate enhanced safety protocols, more effective content filtering, and improved alignment strategies. The demonstrated effectiveness of GAP-generated data for fine-tuning guardrails provides a concrete pathway for improving content moderation systems.
Our assessment indicates that the additional risk introduced by this research is limited, particularly given existing publicly available jailbreaking methods, while the potential benefits for AI safety are substantial. We remain committed to ongoing collaboration with the AI safety community to ensure our research advances robust safeguards while preserving beneficial LLM capabilities.
@@ A.1.1 GAP-Auto
GAP-Auto automates the seed generation process through a two-phase approach. This process involves: (1) Policy Decomposition: High-level content policies are decomposed into specific behavioral constraints using metaprompting techniques with an attacker model (Mistral-123B-v2407), and (2) Seed Generation: For each identified behavior, the system generates both benign and harmful seed prompts, ensuring a balanced dataset. The complete procedure for GAP-Auto seed generation is presented in Algorithm 2.
This aims to solve the initial "cold start" phase, where effective prompts must be generated to initiate the jailbreaking process. Traditionally, this task has been labor-intensive, relying on manually crafted prompts or broad, category-based approaches, which can be time-consuming and limiting. We leverage pre-trained LLMs to dynamically generate diverse seed prompts based on predefined categories or topics, such as those outlined in the OpenAI Safety Usage Policy. By automating the generation of diverse seed prompts, the variant approach not only accelerates the attack generation process but also improves the ability to uncover novel and subtle attack strategies.
This automated approach results in two datasets: GAP-GuardData: A balanced set of benign and harmful prompts derived directly from content policies, and GAP-GuardAttackData: Contains the original benign prompts and the GAP-refined versions of the harmful prompts.
@@
A.1.2 GAP-VLM
Our GAP-VLM variant extends the framework to vision-language models (VLMs) by converting successful text-based jailbreaks into image-embedded attacks. The GAP-VLM pipeline transforms these jailbreak prompts into image + prompt variants specifically designed to circumvent VLM safeguards. The process is formalized in Algorithm 3.
@@ A.2 Performance Analysis
A.2.1 More ASR analysis
In Figure 2, to provide comprehensive insight into GAP’s performance characteristics, we analyze query efficiency from multiple perspectives across all three target models. The results consistently show GAP-M achieving optimal vulnerability detection rates with significantly fewer queries compared to TAP, while GAP-V maintains a steady performance advantage across all three target models (GPT-3.5, Gemma-9B-v2, and Qwen-7B-v2.5).
@@ A.2.1 More ASR analysis
A.2.2 More Metric analysis on Data Quality
In Table 9, we use three complementary metrics to evaluate the diversity of the generated jailbreak prompts.
  • Unique n-grams (%): This metric measures the lexical diversity of the dataset. A higher percentage of unique word sequences indicates less repetitive content and more linguistic variation in the prompts, which is crucial for identifying diverse attack vectors that may not have been previously encountered.
  • Entropy: This metric captures the distributional diversity of the vocabulary. A higher entropy value indicates that the words are more uniformly distributed, and the prompts are more unpredictable, making them more challenging for defenses that rely on a fixed set of keywords or phrases.
  • Self-BLEU: This metric measures the semantic similarity between prompts within the dataset. A lower Self-BLEU score indicates that the prompts are less similar to each other, which confirms that the attack generation process is producing a wide variety of distinct and novel jailbreaks.
GAP-GuardAttackData 94.36 13.72 0.0063
AdvBench seeds chao2023jailbreaking 85.99 8.89 0.1339
JBB seeds chao2024jailbreakbench 81.25 10.27 0.1171
Table 9: Diversity metrics of jailbreak seeds. Higher unique n-grams and entropy indicate greater diversity, while lower Self-BLEU reflects less similarity between prompts. GAP-GuardAttackData outperforms baseline datasets, confirming it generates more linguistically and semantically diverse attacks.
A.3 Implementation Details
Algorithm 1 GAP (Graph of Attacks with Pruning)
1:Query $Q$, branching-factor $b$, maximum width $w$, maximum depth $d$
2:Jailbreak prompt $p$ or failure
[…]
5:  for each leaf node $\ell$ in $G$ do
6:   $C\leftarrow\{\}$ $\triangleright$ Initialize empty set for conversation histories
7:   for each path from root to a leaf in $G$ do
8:     $h\leftarrow$ Concatenate all $[p,r,s]$ tuples in the path
9:     $C\leftarrow C\cup\{h\}$ $\triangleright$ Add path history to set
[…]
12:     $p_{j}\leftarrow\mathcal{A}(Q,global\_context)$ $\triangleright$ Generate prompt using Attacker
13:     $s_{j}\leftarrow$ Retrieve effectiveness of $p_{j}$ based on global_contextglobal\_context
14:   $p_{best}\leftarrow\operatorname*{arg\,max}_{j}s_{j}$
15:   $new\_history\leftarrow\ell.history+[p_{best},\text{response to be generated},\text{score to be calculated}]$
16:   Add child of $\ell$ with prompt $p_{best}$ and history new_historynew\_history
19:  if successful jailbreak found then return jailbreak prompt
21:return failure
Algorithm 2 GAP-Auto Seed Generation
1:High-level content policies
[…]
4:for each behavior $b$ in $B$ do
5:  $s_{benign}\leftarrow$ GenerateBenignPrompt($b$)
6:  $s_{harmful}\leftarrow$ GenerateHarmfulPrompt($b$)
7:  $S_{benign}\leftarrow S_{benign}\cup\{s_{benign}\}$
8:  $S_{harmful}\leftarrow S_{harmful}\cup\{s_{harmful}\}$
9:GAP-GuardData $\leftarrow S_{benign}\cup S_{harmful}$
[…]
11:for each prompt $p$ in $S_{harmful}$ do
12:  $p_{attack}\leftarrow$ ApplyGAP ($p$)
13:  $S_{attack}\leftarrow S_{attack}\cup\{p_{attack}\}$
14:GAP-GuardAttackData $\leftarrow S_{benign}\cup S_{attack}$
15:return GAP-GuardData, GAP-GuardAttackData
Algorithm 3 GAP-VLM Multimodal Attack Generation
1:GAP-generated text prompt $p$
2:Multimodal attack prompt $m$
[…]
8:return $m$
A.3.1 Model Configurations
  • Attacker Models: GAP-M uses Mistral-123B-v2407 while GAP-V uses Vicuna-13B-v1.5